# Moser’s twist theorem revisited

### v2

@@ Abstract.
Area-preserving twist maps, Minimal configurations, Invariant circles
@@ 1. Introduction
The existence of invariant circles (i.e., homotopically non-trivial invariant curves) for area-preserving twist maps was first established by Moser in his seminal work [15], commonly referred to as Moser’s twist theorem. This groundbreaking result is also recognized as one of the foundational contributions to the KAM (Kolmogorov-Arnold-Moser) theory, where the map was initially required to be of class $C^{333}$. The regularity condition was significantly relaxed to $C^{4+\varepsilon}$ by Rüssmann [18]. Subsequently, Moser [16, Page 53] asserted that $C^{3+\varepsilon}$ regularity suffices for the existence of invariant circles by employing more refined estimates within the KAM framework. The full proof of this assertion was independently provided by Rüssmann [19] and Herman [8], each utilizing distinct methodologies. In [19], Rüssmann introduced a novel iteration process that yields invariant circles which are merely continuous. In contrast, Herman’s approach in [8] leverages the Schauder fixed point theorem, building upon his celebrated work on circle diffeomorphisms ([7]). Although Herman’s method is restricted to invariant circles with constant type frequency, it achieves superior smoothness for these circles, specifically $C^{2+\varepsilon}$.
@@
Given $0<\varepsilon\ll 1$, let $V:\mathbb{T}\to\mathbb{R}$ be a $C^{4+\varepsilon}$ function of period 1, where $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$ denotes the flat circle. Let ||$\cdot$|| ⋅ | represent the Euclidean metric in $\mathbb{R}$. The flat metric on $\mathbb{T}$ is defined as
$\|\cdot\|:=\inf_{p\in\mathbb{Z}}|\cdot+p|.$
@@
(1) $G_{n}(x,x^{\prime}):=\frac{1}{2}(x-x^{\prime})^{2}+\frac{1}{q_{n}^{4+\varepsilon}}V(q_{n}x^{\prime}).$
Given $n\in\mathbb{N}$, let $f_{n}:\mathbb{T}\times\mathbb{R}\to\mathbb{T}\times\mathbb{R}$ be the exact area-preserving twist map generated by $G_{n}$.
@@
For each $\alpha\in\mathcal{C}$, there exists a sufficiently large $n\in\mathbb{N}$ such that $f_{n}$ admits an invariant circle with frequency $\alpha$ and the preserved invariant circle is a graph of the function of class $C^{2+\varepsilon^{\prime}}$ for any $\varepsilon^{\prime}<\varepsilon$.
@@
Let $\bar{f}(x,y)=(x+y,y)$. By definition, for each $\varepsilon^{\prime}<\varepsilon$, we have
$\|f_{n}-\bar{f}\|_{C^{3+\varepsilon^{\prime}}}\to 0\quad\text{as}\quad n\to\infty.$
@@
Moser [17] observed that every twist diffeomorphism corresponds to the time-one map of a suitable periodic Hamiltonian system, although the converse does not hold. To ensure the validity of the converse, the condition that there are no conjugate points must also be imposed. Consequently, a result analogous to Theorem 1 can be directly derived for Hamiltonian systems with two degrees of freedom. More precisely, let $\mathbb{T}^{2}$ be the 2-dimensional flat torus. We denote by $\mathcal{D}\subset\mathbb{R}^{2}$ the set of frequency vectors $\omega\in\mathbb{R}^{2}$ satisfying the constant-type Diophantine condition: there exists a constant $D>0$ such that for all non-zero integer vectors $k\in\mathbb{Z}^{2}\setminus\{0\}$,
(2) $|\langle\omega,k\rangle|\geq\frac{D}{\|k\|}.$
@@
$H_{0}(q,p)=\frac{1}{2}\|p\|^{2},\quad p\in\mathbb{R}^{2}.$
@@
The paper is organized as follows. In Section 2, we recall the definition of area-preserving twist maps and discuss relevant properties of minimal configurations with irrational frequencies. Section 3 introduces the definitions of Type I and Type II vectors, along with three criteria for determining the existence of invariant circles. Notably, our subsequent analysis primarily relies on Criterion 3 as this analytical tool. Under the assumptions (19) and (20) stated in Lemma 4.3, Section 4 proves the existence of invariant circles for any given parameter $\kappa$. Furthermore, Section 5 establishes through inductive arguments that the assumptions (19) and (20) are automatically satisfied when $n$ is sufficiently large. For Type I vectors and Criterion 1, our work extends the Katznelson and Ornstein framework [11] while providing more quantitative results compared to their original findings. Consequently, the specially designed Type II vectors and Criterion 3 provide essential foundations for our analysis.
Acknowledgement. The authors would like to thank Zhicheng Tong for his helpful comments. This work is partially under the support of National Natural Science Foundation of China (Grant No. 12122109).
@@ 2. The minimal configuration with frequency $\alpha$
Let us recall some basic notions as preparations. Let $F$ be a diffeomorphism of $\mathbb{R}^{2}$ denoted by $F(x,y)=(X(x,y),Y(x,y))$. Let $F$ satisfy:
  • Periodicity: $F\circ T=T\circ F$ for the translation $T(x,y)=(x+1,y)$;
@@
    $YdX-ydx=dH.$
Then $F$ induces a map on the cylinder denoted by $f$: $\mathbb{T}\times\mathbb{R}\mapsto\mathbb{T}\times\mathbb{R}$. $f$ is called an exact area-preserving twist map. The function $H$: $\mathbb{R}^{2}\rightarrow\mathbb{R}$ is called a generating function of $F$, namely $F$ is generated by the following equations
$\begin{cases}y=-\partial_{1}H(x,x^{\prime}),\\ y^{\prime}=\partial_{2}H(x,x^{\prime}),\end{cases}$
where $F(x,y)=(x^{\prime},y^{\prime})$.
The function $F$ gives rise to a dynamical system whose orbits are given by the images of points of $\mathbb{R}^{2}$ under the successive iterates of $F$. The orbit of the point $(x_{0},y_{0})$ is the bi-infinite sequence
@@
where $(x_{k},y_{k})=F(x_{k-1},y_{k-1})$. The sequence
$(...,x_{-k},...,x_{-1},x_{0},x_{1},...,x_{k},...)$
denoted by ${\bf{x}}:=(x_{i})_{i\in\mathbb{Z}}$ is called a stationary configuration if it satisfies the discrete Euler-Lagrange equation
@@
Let $(x_{i})_{i\in\mathbb{Z}}$ be a minimal configuration with $\alpha$. Then
@@ 3. Criterions on the existence of the invariant circle
Assume $\alpha$ is irrational, then there exists an Aubry-Mather set $\tilde{\mathcal{M}}$ with $\alpha$ (see [2, 6]). For each $z_{0}=(x_{0},y_{0})\in\tilde{\mathcal{M}}$, we have
~=$\tilde{\mathcal{M}}=\overline{\{F^{i}(z_{0})\}}_{i\in\mathbb{Z}},$
where $\bar{A}$ denotes the closure of $A$ in the topology induced by the Riemannian metric on $\mathbb{T}\times\mathbb{R}$. Let $\mathcal{O}$ be an orbit segment in $\tilde{\mathcal{M}}$. Denote $(x_{i},y_{i}):=F^{i}(z_{0})$. The mapping $x_{i}\mapsto i\alpha$ is orientation-preserving from $\mathcal{O}$ to the circle $\mathbb{T}$. A vector $v$ is called a chord if its endpoints are in some orbit segment in $\tilde{\mathcal{M}}$. Denote $\lambda(v)$ the length of the arc between the corresponding multiples of $\alpha$ on $\mathbb{T}$. More precisely, assume the chord $v$ connects $z_{i}=(x_{i},y_{i})$ and $z_{j}=(x_{j},y_{j})$. Then $\lambda(v)=\|(j-i)\alpha\|$. It is clear to see that $\lambda$ is $F$-invariant. It can be viewed as an $F$-invariant distance on $\mathcal{O}$. We also denote
$\theta(v):=x_{j}-x_{i},\quad\Theta(v):=\|\theta(v)\|.$
Based on the orientation-preserving property (see Proposition 2.1), $\lambda(v)$ and $\Theta(v)$ can be related with each other in the sense of average.
@@
Let $v$ be a chord. If $\lambda(v)\geq\frac{2}{q_{N}}$ for $N$ large enough, then
$\left|\frac{1}{q_{N}}\sum_{j=0}^{q_{N}-1}\Theta(F^{j}(v))-\lambda(v)\right|\leq\frac{1}{q_{N}}.$
3.1. $\kappa$-vectors
@@
(6) $10<\varphi(m+1)-\varphi(m)\leq\mathrm{log}_{2}A_{\alpha}+$ 1 .
We also define $\phi(m):=[\varphi(m)]$. We use, once and for all, $[c]$ to denote the largest integer not greater than $c$ for $c\in\mathbb{R}$. It is clear to see that $\phi:\mathbb{N}\to\mathbb{N}$ is nondecreasing. In general, $\phi$ may not be surjective. Meanwhile, $\phi:\mathbb{N}\to\mathbb{N}$ may not be strict increasing. Consequently, $\phi^{-1}:\mathbb{N}\to\mathbb{N}$ may not be well-defined. Denote $S:=\phi(\mathbb{N})$. We introduce $\psi$ as a substitution of $\phi^{-1}:\mathbb{N}\to\mathbb{N}$. For each $\kappa\notin S$, there exists a maximal $\hat{\kappa}\in S$ satisfying $\hat{\kappa}<\kappa$. Note that
@@
$\check{\kappa}:=\left\{\begin{array}[]{ll}\kappa,\quad\kappa\in S,\\ \hat{\kappa},\quad\kappa\notin S.\end{array}\right.$
@@
$\psi(\kappa):=\min\{m\in\mathbb{N}\ |\ \phi(m)=\check{\kappa}\}.$
It follows that $\kappa\to+\infty$ if and only if $\psi(\kappa)\to+\infty$. In the following, we define two types of $\kappa$-vectors. The first one was introduced by [11].
@@
We call $v$ a Type-I $\kappa$-vector, denoted by $v\in(\kappa)_{\text{I}}$, if $\lambda(v)\in[\frac{1}{2^{\kappa+4}},\frac{1}{2^{\kappa}}]$.
To achieve the almost optimal regularity in Theorem 1, we need another kind of $\kappa$-vector.
@@
Assume $v$ connects $(x_{i},y_{i})$ and $(x_{j},y_{j})$. We call $v$ a Type-II $\kappa$-vector, denoted by $v\in(\kappa)_{II}$, if $v\in(\kappa)_{I}$ and
${\frac{1}{(A_{\alpha}+1)^{6}}q_{n_{\kappa}}\leq|i-j|\leq(A_{\alpha}+1)^{6}q_{n_{\kappa}},\quad\lambda(v)\in[\|q_{n_{\kappa}}\alpha\|,2^{4}\|q_{n_{\kappa}}\alpha\|]}$
@@
(9) $\gamma_{0}:=\left[\mathrm{log}_{2}2(A_{\alpha}+2)\right]+1,\quad\bar{N}(\kappa):=n_{\kappa}+2\gamma_{0},$
(10) $\tilde{N}(\kappa):=n_{\kappa}-1,$
where $n_{\kappa}:=\psi(\kappa)$. By (4) and (8), a direct calculation shows that if $v\in(\kappa)_{II}$, then
(11) $\frac{2}{q_{\bar{N}(\kappa)}}<\lambda(v)<\frac{2}{q_{\tilde{N}(\kappa)}}.$
@@
It is easy to verify that the set $(\kappa)_{II}\neq\emptyset$. In fact, the vector $v$ connecting $(x_{i},y_{i})$ and $(x_{i+q_{n_{\kappa}}},y_{i+q_{n_{\kappa}}})$ belongs to $(\kappa)_{II}$.

### v3

@@ Abstract.
Area-preserving twist maps, Minimal configurations, Invariant circles, Type-II chords, Constant type frequencies
@@ 1. Introduction
The existence of invariant circles (i.e., homotopically non-trivial invariant curves) for area-preserving twist maps was first established by Moser in his seminal work [14], commonly referred to as Moser’s twist theorem. This groundbreaking result is also recognized as one of the foundational contributions to the KAM (Kolmogorov-Arnold-Moser) theory, where the map was initially required to be of class $C^{333}$. The regularity condition was significantly relaxed to $C^{4+\varepsilon}$ by Rüssmann [17], where $\varepsilon$ denotes, once and for all, an arbitrarily small positive constant. Subsequently, Moser [15, Page 53] asserted that $C^{3+\varepsilon}$ regularity suffices for the existence of invariant circles by employing more refined estimates within the KAM framework. The full proof of this assertion was independently provided by Rüssmann [18] and Herman [8], each utilizing distinct methodologies. In [18], Rüssmann introduced a novel iteration process that yields invariant circles which are merely continuous. In contrast, Herman’s approach in [8] leverages the Schauder fixed point theorem, building upon his celebrated work on circle diffeomorphisms ([7]). Although Herman’s method is restricted to invariant circles with constant type frequency, it achieves superior smoothness for these circles, specifically $C^{2+\varepsilon}$.
@@
Let $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$ denote the flat circle. Let $|\cdot|$ represent the Euclidean metric in $\mathbb{R}^{d}$. The flat metric on $\mathbb{T}$ is defined as
(1) $\|\cdot\|:=\inf_{p\in\mathbb{Z}}|\cdot+p|.$
@@
Given $k\in[0,+\infty)\cup\{+\infty\}$, let $C^{k}(\mathbb{T})$ denote the space of $C^{k}$ functions on the circle $\mathbb{T}$. Given any $u\in C^{k}(\mathbb{T})$. Let us recall the $C^{k}$-norm:
$\|u\|_{C^{k}}:=\max_{|r|\leq k}\max_{x\in\mathbb{T}}|D^{r}u(x)|,$
where $r:=(r_{1},\ldots,r_{n})$, $|r|:=r_{1}+\cdots+r_{n}$.
@@
(2) $G_{n}(x,x^{\prime}):=\frac{1}{2}(x-x^{\prime})^{2}+\frac{1}{q_{n}^{4+\varepsilon}}V(q_{n}x^{\prime}),$
where $V\in C^{4+\varepsilon}(\mathbb{T})$ and $q_{n}$ denotes the denominator of the $n$-th convergent of $\alpha$. This means that the construction of $G_{n}$ depends on $\alpha$; for notational simplicity, we suppress the subscript $\alpha$ in our notation. Given $n\in\mathbb{N}$, let $F_{n}:\mathbb{T}\times\mathbb{R}\to\mathbb{T}\times\mathbb{R}$ be the exact area-preserving twist map generated by $G_{n}$.
@@
For each $\alpha\in\mathcal{C}$, there exists a sufficiently large $n\in\mathbb{N}$ such that $F_{n}$ admits an invariant circle with frequency $\alpha$ and the preserved invariant circle can be represented as a graph of the function of class $C^{2+\varepsilon^{\prime}}$ for any $\varepsilon^{\prime}<\varepsilon$.
@@
Let $T(x,y)=(x+y,y)$. By definition, for each $\varepsilon^{\prime}<\varepsilon$, we have
$\|F_{n}-T\|_{C^{3+\varepsilon^{\prime}}}\to 0\quad\text{as}\quad n\to\infty.$
@@
Moser [16] observed that every $C^{\infty}$ twist diffeomorphism corresponds to the time-one map of a suitable $C^{\infty}$ periodic Hamiltonian system, although the converse does not hold. To ensure the validity of the converse, the condition that there are no conjugate points must also be imposed. Consequently, a result analogous to Theorem 1 can be directly derived for Hamiltonian systems with two degrees of freedom. More precisely, let $\mathbb{T}^{2}$ be the 2-dimensional flat torus. We denote by $\mathcal{D}\subset\mathbb{R}^{2}$ the set of frequency vectors $\omega\in\mathbb{R}^{2}$ satisfying the constant-type Diophantine condition: there exists a constant $D>0$ such that for all non-zero integer vectors $k\in\mathbb{Z}^{2}\setminus\{0\}$,
(3) $|\langle\omega,k\rangle|\geq\frac{D}{|k|}.$
@@
$H_{0}(q,p)=\frac{1}{2}|p|^{2},\quad p\in\mathbb{R}^{2}.$
@@
Strategy of the proof of Theorem 1. Katznelson and Ornstein [11] introduced a criterion for the existence of invariant circles (referred to as Criterion 1 in this paper), expressed in terms of certain vectors with endpoints on orbit segments (called Type-I chords here; see Definition 3.3). However, working directly with such chords makes it difficult to achieve almost optimal result on the regularity of the perturbation for twist maps. We therefore introduce a more refined class of chords (called Type-II chords; see Definition 3.4), whose definition imposes an additional constraint on the “difference of positions” of their endpoints on the minimal configuration.
Based on these Type-II chords, we formulate a new criterion for invariant circles (Criterion 2) and prove its equivalence to Criterion 1 (Lemma 3.6). To connect the construction with circle diffeomorphism theory, we further reduce Criterion 2 to Criterion 3, which is a boundedness condition on $\widetilde{\mathbf{K}}^{0}$ (defined in (17)) related to a certain multiplicative cocycle. This reduction is carried out in Section 3.
In light of the idea to prove the Denjoy inequality for circle diffeomorphisms, Katznelson and Ornstein [11] established a corresponding result for twist maps (Proposition 4.1), which implies the quantity $\widetilde{\mathbf{K}}^{0}$ can be controlled by a certain distortion $\widetilde{\mathbf{K}}^{1}$ (defined in (22)). Introducing the second and third-order difference quotients $\nabla^{1}$ and $\nabla^{2}$ of the minimal configuration (see definitions (23) and (24)), and under suitable hypotheses on $\nabla^{1},\nabla^{2}$ (conditions (27) and (28)), one obtains exponential decay of $\widetilde{\mathbf{K}}^{1}$ (Lemma 4.3). Boundedness of $\widetilde{\mathbf{K}}^{0}$ then follows from Proposition 4.1.
The proof of Lemma 4.3 is most delicate part in this paper. Establishing the exponential decay of $\widetilde{\mathbf{K}}^{1}$ requires several steps. We first prove a monotonicity property for $\nabla^{1}$ (Lemma 4.4). When the orbit segment is short relative to the scale of the chord-pair involved in $\nabla^{1}$, we show exponential decay of $\widetilde{\mathbf{K}}^{1}$ (Lemma 4.6). Using the assumption on $\nabla^{2}$, we then extend this decay to the case where the length of the orbit segment matches the scale involved in $\nabla^{1}$ (Lemma 4.8). Finally we obtain, a posteriori, a sharpened exponential decay estimate for $\widetilde{\mathbf{K}}^{1}$ (Proposition 4.9). These arguments are contained in Section 4.
The last step consists of an induction argument showing that there exists a scale $\kappa_{0}$ such that, on all scales finer than $\kappa_{0}$, the conditions on $\nabla^{1}$ and $\nabla^{2}$ ((27) and (28)) are automatically satisfied (Lemma 5.1). The initial scale $\kappa_{0}$ depends in a precise way on the size of the perturbation of the system (Lemma 3.7). This is presented in Section 5, together with a basic estimate for the vertical-coordinate difference between consecutive points of the orbit (Lemma 2.2) given in Section 2. To deduce the final result from the induction hypothesis, we need to control the growth of $\nabla^{1}$ in relation to $\nabla^{2}$ (Lemma 5.4) as the scale of the chord increases, and then combine this with the refined estimate for $\nabla^{2}$ (see (63)) obtained by Katznelson and Ornstein ([11, Corollary 3.1]) to complete the proof of Theorem 1.
Roughly speaking, the main idea to prove Theorem 1 is to prove that an Aubry-Mather set which “looks like” an invariant circle must indeed be an invariant circle. The precise meaning of “looks like” is encoded by the second and third-order difference quotients $\nabla^{1}$ and $\nabla^{2}$ of the minimal configuration. Through a bootstrap-type procedure, such an Aubry-Mather set cannot contain any “gaps” and therefore necessarily satisfies Criterion 1.
Acknowledgement. The authors would like to thank Zhicheng Tong for his helpful comments. This work is partially under the support of National Natural Science Foundation of China (Grant No. 12122109) and the Beijing Natural Science Foundation (Grant No. QY25241).
@@ 2. The minimal configuration with frequency $\alpha$
Let us recall some basic notions as preparations. Let $\tilde{F}$ be a diffeomorphism of $\mathbb{R}^{2}$ denoted by $\tilde{F}(x,y)=(X(x,y),Y(x,y))$. Let $\tilde{F}$ satisfy:
  • Lift condition: $\tilde{F}$ is isotopic to the identity;
@@
    YdXydx=dH.YdX-ydx=dH.
Then $\tilde{F}$ induces a map on the cylinder denoted by $f$: $\mathbb{T}\times\mathbb{R}\mapsto\mathbb{T}\times\mathbb{R}$. $f$ is called an exact area-preserving twist map. The function $H$: $\mathbb{R}^{2}\rightarrow\mathbb{R}$ is called a generating function of $\tilde{F}$, namely $\tilde{F}$ is generated by the following equations
$\begin{cases}y=-\partial_{1}H(x,X),\\ Y=\partial_{2}H(x,X),\end{cases}$
where $\tilde{F}(x,y)=(X,Y)$.
The function $\tilde{F}$ gives rise to a dynamical system whose orbits are given by the images of points of $\mathbb{R}^{2}$ under the successive iterates of $\tilde{F}$. The orbit of the point $(x_{0},y_{0})$ is the bi-infinite sequence
@@
where $(x_{k},y_{k})=\tilde{F}(x_{k-1},y_{k-1})$. The sequence
[…]
denoted by ${\bf{x}}:=(x_{i})_{i\in\mathbb{Z}}$ is called a stationary configuration associated to $\tilde{F}$ if it satisfies the discrete Euler-Lagrange equation
@@
Let $(x_{i})_{i\in\mathbb{Z}}$ be a minimal configuration with rotation number $\alpha$. Then
@@ 3. Criterions on the existence of the invariant circle
Let $F:\mathbb{T}\times\mathbb{R}\to\mathbb{T}\times\mathbb{R}$ be an exact area-preserving twist map defined on $\mathbb{T}\times\mathbb{R}$. Take an irrational number $\alpha$. It is known (see [2, 6]) that there exists an Aubry–Mather set $\widetilde{\mathcal{M}}_{\alpha}\subset\mathbb{T}\times\mathbb{R}$.
Let $\pi:\mathbb{T}\times\mathbb{R}\to\mathbb{T}$ be the standard projection. Briefly, the projected set $\pi\widetilde{\mathcal{M}}$ onto $\mathbb{T}$ consists of all minimal configurations with rotation number $\alpha$. For simplicity, we do not distinguish here between a minimal configuration on $\mathbb{T}$ and its lift to $\mathbb{R}$. Denote by $H:\mathbb{R}^{2}\to\mathbb{R}$ the generating function of the lift $\tilde{F}$. More precisely, let $\mathbf{x}=(x_{i})_{i\in\mathbb{Z}}$ be a minimal configuration generated by $H$. Define
$y_{0}:=-\partial_{1}H(x_{0},x_{1}).$
Let
$\mathcal{O}:=\{F^{i}(x_{0},y_{0})\}_{i\in\mathbb{Z}},$
where $\mathcal{O}$ is a complete orbit passing through $(x_{0},y_{0})$. If the rotation number $\rho(\mathbf{x})$ is irrational, and we set $\alpha=\rho(\mathbf{x})$, then $\widetilde{\mathcal{M}}_{\alpha}$ can be expressed as
$\widetilde{\mathcal{M}}_{\alpha}=\overline{\mathcal{O}},$
where $\overline{\mathcal{O}}$ denotes the closure of $\mathcal{O}$ in the topology induced by the Riemannian metric on $\mathbb{T}\times\mathbb{R}$.
The mapping $x_{i}\mapsto i\alpha$ is orientation-preserving from $\mathcal{O}$ to the circle $\mathbb{T}$. A vector $v$ is called a chord if its endpoints lie on some orbit segment of $\mathcal{O}$. Denote by $\lambda_{1}(v)$ the length of the arc between the corresponding multiples of $\alpha$ on $\mathbb{T}$. Concretely, if the chord $v$ connects $z_{i}=(x_{i},y_{i})$ and $z_{j}=(x_{j},y_{j})$, then
$\lambda_{1}(v)=\|(j-i)\alpha\|,$
where $\|\cdot\|$ is defined by (1). It is clear that $\lambda$ is invariant under $F$. For clarity, $\lambda_{1}(v)$ is called the $\lambda$-length of $v$ on $\mathcal{O}$. Under the same assumption that the chord $v$ connects $z_{i}=(x_{i},y_{i})$ and $z_{j}=(x_{j},y_{j})$, we also write
$\theta(v):=x_{j}-x_{i},\qquad\Theta_{1}(v):=\|\theta(v)\|.$
The quantity $\Theta_{1}(v)$ is called the $\vartheta$-length of $v$ on $\mathcal{O}$.
Remark 3.1.
By the definitions of the $\lambda$-length and the $\vartheta$-length of $v$ on $\mathcal{O}$, for two chords $v_{1}$ and $v_{2}$ on $\mathcal{O}$, even if their $\lambda$-distances coincide, their $\vartheta$-lengths may still differ. In $\widetilde{\mathcal{M}}_{\alpha}$ we fix, once and for all, an arbitrary complete orbit $\mathcal{O}$. All subsequent discussions about a chord $v$ refer to the vectors that arise from some orbit segment whose endpoints belong to this chosen complete orbit $\mathcal{O}$.
Based on the orientation-preserving property (see Proposition 2.1), $\lambda_{1}(v)$ and $\Theta_{1}(v)$ can be related with each other in the sense of average. Recall $q_{n}$ denotes the denominator of the $n$-th convergent of $\alpha$.
@@
Let $v$ be a chord. If $\lambda_{1}(v)\geq\frac{2}{q_{N}}$ for $N$ large enough, then
$\left|\frac{1}{q_{N}}\sum_{j=0}^{q_{N}-1}\Theta_{1}(F^{j}(v))-\lambda_{1}(v)\right|\leq\frac{1}{q_{N}}.$
3.1. $\kappa$-chords
@@
(7) $0<\varphi(m+1)-\varphi(m)\leq\mathrm{log}_{2}A_{\alpha}+1.$
We also define
(8) $\phi(m):=[\varphi(m)].$
We use, once and for all, $[c]$ to denote the largest integer not greater than $c$ for $c\in\mathbb{R}$. It is clear to see that $\phi:\mathbb{N}\to\mathbb{N}$ is nondecreasing. In general, $\phi$ may not be surjective. Meanwhile, $\phi:\mathbb{N}\to\mathbb{N}$ may not be strict increasing. Consequently, $\phi^{-1}:\mathbb{N}\to\mathbb{N}$ may not be well-defined. Denote $S:=\phi(\mathbb{N})$. We introduce $\psi$ as a substitution of $\phi^{-1}:\mathbb{N}\to\mathbb{N}$. For each $\kappa\notin S$, there exists a maximal $\hat{\kappa}\in S$ satisfying $\hat{\kappa}<\kappa$. Note that
@@
$\check{\kappa}:=\left\{\begin{array}[]{ll}\hskip-3.99994pt\kappa,\quad\kappa\in S,\\ \hskip-3.99994pt\hat{\kappa},\quad\kappa\notin S.\end{array}\right$.
@@
(10) $\psi(\kappa):=\min\{m\in\mathbb{N}\ |\ \phi(m)=\check{\kappa}\}.$
It follows that $\kappa\to+\infty$ if and only if $\psi(\kappa)\to+\infty$. In the following, we define two types of $\kappa$-chords. The first one was introduced by [11].
@@
We call $v$ a Type-I $\kappa$-chord, denoted by $v\in(\kappa)_{\text{I}}$, if $\lambda_{1}(v)\in[\frac{1}{2^{\kappa+4}},\frac{1}{2^{\kappa}}]$.
To achieve the almost optimal regularity in Theorem 1, we need another kind of $\kappa$-chord.
@@
Assume $v$ connects $(x_{i},y_{i})$ and $(x_{j},y_{j})$. We call $v$ a Type-II $\kappa$-chord, denoted by $v\in(\kappa)_{II}$, if $v\in(\kappa)_{I}$ and
$\frac{1}{(A_{\alpha}+1)^{6}}q_{n_{\kappa}}\leq|i-j|\leq(A_{\alpha}+1)^{6}q_{n_{\kappa}},\quad\lambda_{1}(v)\in[\|q_{n_{\kappa}}\alpha\|,2^{4}\|q_{n_{\kappa}}\alpha\|]$}
@@
(12) $\gamma_{0}:=\left[\mathrm{log}_{2}2(A_{\alpha}+2)\right]+4,\quad\bar{N}(\kappa):=n_{\kappa}+2\gamma_{0},$
(13) $\tilde{N}(\kappa):=n_{\kappa}-2\gamma_{0},$
where $n_{\kappa}=\psi(\kappa)$. By (5), (11), (12) and (13), a direct calculation shows that if $v\in(\kappa)_{II}$, then
(14) $\frac{2}{q_{\bar{N}(\kappa)}}<\lambda_{1}(v)<\frac{1}{4q_{\tilde{N}(\kappa)}}.$
@@
  • It is easy to verify that the set $(\kappa)_{II}$ is nonempty. In fact, the chord $v$ joining $(x_{i},y_{i})$ and $(x_{i+q_{n_{\kappa}}},y_{i+q_{n_{\kappa}}})$ belongs to $(\kappa)_{II}$.
  • We illustrate the notations with two explicit constant-type irrational numbers.
    • For $\alpha$ equal to the golden ratio $\frac{\sqrt{5}-1}{2}=[0;1,1,\dots]$, we have $A_{\alpha}=1$ and $\gamma_{0}=\bigl[\log_{2}\bigl(2(A_{\alpha}+2)\bigr)\bigr]+4=6$. Because $0<\varphi(m+1)-\varphi(m)\leq 1$, the map $\phi$ is surjective: $\phi(\mathbb{N})=\mathbb{N}$. Simple computations give $\psi(1)=1,\;\psi(5)=9,\;\psi(10)=17,\;\psi(100)=146,\;\psi(500)=722$.
    • For $\alpha=\sqrt{2}-1=[0;2,2,\dots]$, we have $A_{\alpha}=2$ and $\gamma_{0}=\bigl[\log_{2}\bigl(2(A_{\alpha}+2)\bigr)\bigr]+4=7$. Here $\phi$ is not surjective; indeed $\phi(\mathbb{N})=\mathbb{N}\setminus\mathcal{T}$ with $\mathcal{T}=\{4,9,14,18,23,28,\dots\}$, where the gap between two consecutive elements of $\mathcal{T}$ is either $4$ or $5$. Calculations yield $\psi(1)=1,\;\psi(4)=\psi(3)=4,\;\psi(28)=23,\;\psi(100)=80,\;\psi(500)=395$.